FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE A^{2}Π–X^{2}Σ^{+} (RED) SYSTEM OF ^{13}C^{14}N.

\underline{R.S. RAM}, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK; Department of Chemistry, University of Arizona, Tucson, AZ 85721, USA; and P.F. BERNATH, Department of Chemistry and Biochemistry, Old Dominion University, Norfolk, VA 23529 USA; Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK.

High resolution emission spectra of the A^{2}Π–X^{2}Σ^{+} transition of ^{13}C^{14}N have been measured in the 15000–24000 cm^{−1} region. Molecules were produced by the reaction of ^{13}CH_{4} and ^{14}N_{2} in an active nitrogen afterglow discharge. The spectra were recorded using the Fourier transform spectrometer associated with the McMath-Pierce Solar Telescope of the National Solar Observatory. A rotational analysis of 27 bands involving the excited state vibrational levels v′ = 9–22 and the ground state vibrational levels up to v′′ = 12 has been obtained. A much improved set of spectroscopic constants have been determined for the v = 0–22 vibrational levels of the A^{2}Π state by combining the present measurements with those reported previously for the v = 0–8 vibrational levels of the A^{2}Π state^{a} and existing infrared and millimeter-wave measurements of ^{13}C^{14}N. The 6–3, 7–4, 8–5 and 9–6 bands of the B^{2}Σ^{+}–A^{2}Π transition were also identified in the 23300– 24000 cm^{−1} region and were included in the final analysis. An experimental line list and calculated term values are provided.

aRam et al., Astrophys. J. Suppl. Ser., 188, (2010), 500.